### v2

# TUN: Detecting Significant Points in Persistence Diagrams with Deep Learning

Abstract
Persistence diagrams (PDs) provide a powerful tool for understanding the topology of the underlying shape of a point cloud. However, identifying which points in PDs encode genuine signals remains challenging. This challenge directly hinders the practical adoption of topological data analysis in many applications, where automated and reliable interpretation of persistence diagrams is essential for downstream decision-making. In this paper, we study automatic significance detection for one-dimensional persistence diagrams. Specifically, we propose TopologyUnderstandingNet (TUN), a multi-modal network that combines enhanced PD descriptors with self-attention, a PointNet-style point cloud encoder, learned fusion, and per-point classification, alongside stable preprocessing and imbalance-aware training. It provides an automated and effective solution for identifying significant points in PDs, which are critical for downstream applications. Experiments show that TUN outperforms classic methods in detecting significant points in PDs, illustrating its effectiveness in real-world applications.
I Introduction
Topological Data Analysis (TDA), particularly through the framework of persistent homology, has emerged as a powerful mathematical approach for extracting topological information from complex data. By constructing multi-scale representations of data, persistent homology robustly quantifies essential topological features such as connected components, loops, and voids, summarizing them in concise descriptors known as persistence diagrams (PDs) [13]. A point in a PD represents the birth time and death time of a topological feature.
A central challenge within TDA is achieving accurate topology understanding, particularly for widely used point cloud representations. Point clouds have become a fundamental representation of three-dimensional objects and environments, with widespread applications across scientific and engineering disciplines, including robotics [44], medical imaging [34], and computational biology [42]. Topology understanding is crucial for processing geometric models represented by point clouds, as the topology can have a decisive impact on the model’s functionality. When a point cloud is sampled from a geometric shape, its PD holds the potential to reveal intrinsic topological invariants (for example, the genus, the Betti number, and the loop or void features), providing insights into the fundamental structure of the original object. Accurate topology understanding can help us perform more reliable analysis or performance prediction of the model.
Figure 1: An illustration of identifying significant points in 1-PDs. (a)(b): point clouds sampled from two CAD models with genus 2 and 0, respectively. (c)(d): corresponding 1-PDs computed by generating alpha filtrations. (e)(f): identify significant points in 1-PD.
However, the promise of automated topological discovery from persistence diagrams remains largely unfulfilled. The interpretation of PDs constitutes a challenging bottleneck in the TDA pipeline. A PD is a multiset of birth-death pairs, and the core problem involves distinguishing points corresponding to genuine, significant topological features from those representing topological noise or minor geometric details. Traditional heuristics often assume that points with high persistence (long lifespan) are significant [14], while those near the diagonal represent noise. Yet, this simple thresholding approach proves unsatisfactory in practice [6], as the concept of “significance” is inherently context-dependent. Accurate interpretation requires comprehensive assessment, considering not only persistence but also absolute birth and death times, relationships to the global distribution of other points in the diagram, and crucially, the geometric properties of the original point cloud. This complex analysis has traditionally relied on manual expertise, demanding a deep understanding of algebraic topology and rendering the process subjective, time-consuming, and difficult to scale.
The primary goal of this study is to interpret one-dimensional persistence diagrams (1-PDs) computed by alpha filtration [13] so as to recover the genuine 1-d loop features of the underlying geometric shapes from which the point clouds are sampled. These loop features can be essential for downstream tasks such as property prediction and control. Accordingly, a point in a 1-PD is deemed significant only if it corresponds to such a true 1-d loop of the original model. Because the topology of these shapes can be intricate, the decision must exploit both the internal structure of the persistence diagram and the geometric information carried by the point cloud itself. Classic persistence-threshold methods [14, 17], which usually ignore the link between diagram features and the topological invariants of the shape, are insufficient for this purpose.
Figure 1 presents concrete examples that illustrate the challenges of automatic and accurate identification of significant points in persistence diagrams. For the point cloud in Figure 1(a), sampled from a surface with genus 2 (thus containing 4 significant loop features), the corresponding 1-PD is shown in Figure 1(c). Due to dense sampling, these significant loop features appear at early birth times while maintaining relatively large persistence values. Figure 1(e) highlights the points corresponding to these significant loop features. Notably, although another point exhibits relatively large persistence, its late birth time indicates that it does not represent any significant loop feature of the original sampled surface and should not be classified as significant. Similarly, the point cloud in Figure 1(b), sampled from a genus 0 surface (containing no significant loop features), produces a one-dimensional persistence diagram with no significant points, as shown in Figure 1(d) and (f). However, this diagram still contains a point with relatively large persistence but a late birth time. Moreover, even visually similar persistence diagrams may yield different sets of significant points when the underlying point clouds possess distinct geometric characteristics. Consider the pair of 1-PDs shown in Figure 2(a) and (b): both are densely sampled from planar shapes, yet the first is sampled from two adjacent ellipses while the second contains only a similar single closed curve. Although the resulting diagrams are nearly indistinguishable, 1-PD in (a) exhibits two significant points, whereas 1-PD in (b) contains only one. Crucially, the second-most persistent point in (b) is born later than its counterpart in (a), confirming that the geometric context encoded in the point cloud directly governs the topological interpretation of the persistence diagram. These examples clearly demonstrate that relying solely on persistence as a criterion for identifying significant points cannot provide an accurate topological understanding of the underlying geometric models represented by point clouds.
Figure 2: Examples illustrating the challenges of identifying significant points in 1-PDs. (a)(b): two 1-PDs derived from similar planar shapes, yet with distinct sets of significant points.
To address these challenges, we present Topology Understanding Net (TUN), a deep learning framework that formulates topology understanding of PD as per‑point classification: each diagram point is labeled as significant or noise. TUN uses a multi‑modal architecture that fuses topological descriptors from persistence diagrams with geometric context from the raw point cloud, enabling the model to resolve ambiguities that PD‑only heuristics cannot by learning how geometry and topology jointly determine feature significance. In this study, we focus on one‑dimensional persistence diagrams (1‑PDs), where identifying significant points is essential for understanding loop patterns in curves and surfaces. In our setting, 1‑PDs are computed from the alpha filtration [13]; significant points correspond to generators of the first homology group, and their count equals the first Betti number. The formulation can naturally extend to higher‑dimensional diagrams. However, a major obstacle for this task has been the absence of datasets for supervised learning. We therefore construct the first large‑scale dataset for topology understanding, providing ground‑truth labels for significant points in computed 1‑PDs across diverse categories of point clouds. Experiments show that TUN consistently outperforms traditional methods and attains near‑perfect accuracy, offering a practical tool that bridges abstract topological representations with concrete geometric data for large‑scale analysis.
In summary, our main contributions are: (1) We construct and release the first large-scale, manually labeled dataset specifically designed for supervised learning of topological feature significance in 1-PDs. (2) We propose Topology Understanding Net (TUN), the first deep learning framework for automated topology understanding of persistence diagrams, which uniquely leverages both topological and geometric data through a multi-modal architecture. (3) Experimental results demonstrate that TUN achieves state-of-the-art performance, significantly outperforming traditional methods across four diverse categories of datasets.
II Related Work
In this section, we review related work concerning topological data analysis and topology understanding of PD.
II-A Topological Data Analysis (TDA)
Topological Data Analysis  [12, 13, 46, 38] offers mathematically grounded tools to study the shape of data in a way that is robust to noise and invariant to many sampling artifacts. The central construct is persistent homology, which tracks the birth and death of topological features across a filtration and summarizes them in persistence diagrams (PDs). Foundational texts and surveys have established theoretical properties (such as the computational methods of PDs for different kinds of filtrations [13, 27] and the stability of PDs under perturbations [10]) and have documented successful applications across disciplines, such as molecular science [42, 3, 29, 41], materials science [40, 20, 25, 30, 35], cognitive science [26, 8], geometric design and processing [5, 11, 43, 17, 18, 9, 15], and medical image processing [4, 34].
In particular, TDA has been integrated with machine learning frameworks to enhance feature extraction and classification tasks [31, 2]. For example, PDs can be integrated into a deep learning framework as a topological prior. PersLay [7] embeds PDs into a vector space using learnable functions, enabling them to be processed by standard neural network layers. Persformer [33] embeds persistence diagrams into deep learning by treating each diagram as an unordered set of points and utilizing a Transformer to learn hidden features without hand-crafted vectorization. But these frameworks cannot be used to give point-wise significance prediction directly. Moreover, many studies integrate PDs into loss functions to give more topology-aware training results [39, 21, 22]
In summary, these applications usually position PDs as robust and expressive descriptors for complex data across various fields. Notice that a lot of studies use point clouds as the input data to compute PDs, and then use PDs as the feature representation for downstream tasks. Therefore, the ability to accurately understand the topological features in PDs is essectial.
II-B Topology understanding of PD
Understanding the significance of points in PDs has been explored using statistical and geometric approaches. A common heuristic ranks points by persistence and declares the top $k$ (or the single most persistent point) as significant, discarding the remainder as noise [5, 11]. A more principled statistical alternative constructs a confidence set around the diagonal of a PD via bootstrap resampling to separate signal from noise [14]. This procedure depends on the underlying point-cloud density and requires user-selected parameters, limiting automation. Clustering-based strategies (e.g., $k$-means with $k=2$) similarly partition PD points into significant and noise classes [17, 9], but for cases that a PD contains only significant points or only noise points, this method is not feasible. In summary, these families of methods ultimately rely on persistence as the primary importance measure, which is intrinsically limited [6]: persistence may omit the geometric context that generated the diagram. Low-persistence features may be genuinely meaningful when small-scale features are present, whereas high-persistence features may be artifacts of outliers.
III Preliminaries
In this section, we review the foundational concepts related to our methodology, including alpha filtration, persistence homology, and persistence diagram.
III-1 Alpha Filtration
The convex hull of $n+1$ affinely independent points $\{u_{0},u_{1},\ldots,u_{n}\}\subset\mathbb{R}^{N}$ constitutes an $n$-simplex, denoted $u_{0},u_{1},\ldots,u_{n}]$. Foundational geometric primitives such as vertices, edges, triangles, and tetrahedra correspond to 0-, 1-, 2-, and 3-simplices, respectively. A simplicial complex $K$ is a finite set of simplices where two conditions hold: any face of a simplex in $K$ must also belong to $K$, and the intersection of any two simplices in $K$ is either empty or a shared face [13]. The dimensionality of $K$ is determined by the highest dimension of any simplex it contains. A simplicial complex $L$ is a subcomplex of $K$ if $L\subseteq K$.
The alpha complex is a specific type of simplicial complex constructed from a point cloud, and its definition relies on the concepts of Voronoi and Delaunay complexes. Given a point cloud $P=\{x_{i}\}_{i=1}^{m}\subset\mathbb{R}^{n}$, the Voronoi cell for a point $x_{i}$ is the region of space closer to $x_{i}$ than to any other point in $P$:
$V_{x_{i}}=\{x\in\mathbb{R}^{n}:\|x-x_{i}\|\leq\|x-x_{j}\|,\forall j\neq i\},$
where $\|\cdot\|$ is the Euclidean norm. The dual graph of the Voronoi tessellation gives the Delaunay complex, which includes a simplex for every subset of points whose Voronoi cells have a non-empty common intersection:
$\operatorname{Del}(P)=\left\{[x_{i_{1}},\cdots,x_{i_{q}}]:\bigcap_{j=1}^{q}V(x_{i_{j}})\neq\emptyset\right\}.$
The alpha complex is a subset of the Delaunay complex, controlled by a scale parameter $r\geq 0$. For each point $x\in P$, consider the intersection of its Voronoi cell $V_{x}$ with a closed ball of radius $r$ centered at $x$, denoted $R_{x}(r)=B_{x}(r)\cap V_{x}$. The alpha complex $\operatorname{Alp}(P,r)$ is then formed by all simplices whose constituent points have intersecting restricted Voronoi regions:
$\operatorname{Alp}(P,r)=\left\{[x_{i_{1}},\dots,x_{i_{q}}]:\bigcap_{j=1}^{q}R_{x_{i_{j}}}(r)\neq\emptyset\right\}.$
The condition $R_{x}(r)\subseteq V_{x}$ ensures that $\operatorname{Alp}(P,r)$ is always a subcomplex of $\operatorname{Del}(P)$.
By systematically increasing the radius $r$ from zero, we generate a nested sequence of alpha complexes. This ordered family of complexes,
$\emptyset=\operatorname{Alp}(P,r_{0})\subset\operatorname{Alp}(P,r_{1})\subset\cdots\subset\operatorname{Alp}(P,r_{k})\subset\cdots,$
where $0\leq r_{0}<r_{1}<\dots$, is known as the alpha filtration of the point cloud $P$. Figure 3 provides a visual depiction of this process.
III-2 Persistent Homology and Persistence Diagrams
A powerful method in computational topology, persistent homology analyzes the evolution of topological features—such as connected components, tunnels, and voids—across a sequence of nested spaces [12, 13]. It quantifies the lifespan of these features, distinguishing robust structures from ephemeral noise.
The formalism begins with the algebraic structure of a simplicial complex $K$. For a given dimension $n$, the set of all $n$-simplices in $K$ forms a basis for the $n$-chain group, $C_{n}(K)$. An element of this group, an $n$-chain, is a formal sum $c=\sum a_{i}\sigma_{i}$, where each $\sigma_{i}$ is an $n$-simplex and the coefficients $a_{i}$ are taken from a field, typically $\mathbb{Z}_{2}$ in computational settings. The boundary operator $\partial_{n}:C_{n}(K)\to C_{n-1}(K)$ maps an $n$-simplex $\sigma=[x_{0},x_{1},\ldots,x_{n}]$ to the sum of its (n1)(n-1)-dimensional faces:
$\partial_{n}\sigma=\sum_{j=0}^{n}[x_{0},\ldots,\hat{x}_{j},\ldots,x_{n}],$
where $\hat{x}_{j}$ indicates that vertex $x_{j}$ is omitted. This operator allows us to define two important subgroups of $C_{n}(K)$: the group of $n$-cycles, $Z_{n}(K)=\operatorname{Ker}(\partial_{n})$, which are chains with no boundary, and the group of $n$-boundaries, $B_{n}(K)=\operatorname{Im}(\partial_{n+1})$, which are chains that are boundaries of $(n+1)$-chains. The $n$-th homology group is the resulting quotient group $H_{n}(K)=Z_{n}(K)/B_{n}(K)$, whose elements represent the $n$-dimensional holes of the complex.
Figure 3: An example of alpha filtration. As the filtration parameter increases, more and more simplices appear.
Figure 4: Topological interpretation of PD. (a) A planar point cloud with five significant loop features. (b) Its 1-PD. (c) The topology understanding of the confidence set. (d) The topology understanding of the 2-means clustering.
Figure 5: The entire network architecture of TUN.
When applied to a filtration, such as the alpha filtration described previously,
$K_{0}\subset K_{1}\subset\cdots\subset K_{m},$
the inclusion map $K_{i}\hookrightarrow K_{j}$ for $i\leq j$ induces a homomorphism on the corresponding homology groups, $f_{n}^{i,j}:H_{n}(K_{i})\to H_{n}(K_{j})$. The $n$-th persistent homology group from $i$ to $j$ is the image of this homomorphism, $H_{n}^{i,j}=\operatorname{Im}f_{n}^{i,j}$. Its rank, $\beta_{n}^{i,j}=\operatorname{rank}H_{n}^{i,j}$, is the $n$-th persistent Betti number. Persistent homology thus tracks which homology classes (features) “persist” as the filtration progresses, noting when they are born and when they die (i.e., merge into an older class).
The output of persistent homology is commonly summarized in a persistence diagram (PD), a multiset of points in the extended plane $\mathbb{R}^{2}\cup\{\infty\}$ [13]. For each dimension $n$, the $n$-th PD contains a point $(b_{i},d_{i})$ for each $n$-dimensional topological feature that appears at filtration value $b_{i}$ (its birth time) and disappears at value $d_{i}$ (its death time). The vertical distance from a point to the diagonal, $d_{i}-b_{i}$, is its persistence, measuring the feature’s lifespan within the filtration.
Based on their persistence, points in a PD are often classified into two types. Those far from the diagonal, indicating high persistence, are considered significant points and correspond to robust topological features (significant $n$-cycles). Conversely, points near the diagonal have low persistence and are typically interpreted as noise points, representing transient or trivial features (noise $n$-cycles).
III-3 Topology understanding of PD
Now we briefly review the confidence-set method and the 2-means clustering method for topology understanding of PDs. Figure 4(a) depicts a planar point cloud whose underlying shape contains five significant loops; (b) displays the corresponding 1-PD, in which five points exhibit comparatively large persistence. For the confidence set approach (c), a statistically derived band (red) is drawn around the diagonal (with confidence level 0.5), and any point falling inside this band is regarded as noise, whereas points outside are deemed significant [14]. In (d), all PD points are orthogonally projected onto the anti-diagonal line y=xy=-x and then partitioned into two clusters via 2-means. The cluster farther from the diagonal is taken as the significant set, and the closer cluster as noise.
IV Methods
In this section, we introduce our novel deep learning framework, the TUN, designed for the automated classification of significance in persistence diagrams (PDs). The model leverages a multi-modal architecture that jointly processes persistence diagrams and their corresponding raw point cloud data to deliver robust and accurate predictions. Our TUN network consists of four main components: a Persistence Diagram Encoder, a Point Cloud Encoder, a Multi-Modal Fusion module, and a final Significance Classifier. The overall architecture is depicted in Figure 5. More details of implementation and hyperparameters of our TUN model are available in Appendix -A.
IV-A Input Feature Extraction and Processing
Our model processes three distinct types of input data: the persistence diagram, the raw point cloud, and a set of auxiliary features.
Persistence Diagram Features.
The primary topological input is a one-dimensional persistence diagram, represented as a set of 2D points $P=\{p_{i}=(b_{i},d_{i})\}_{i=1}^{N}$, where $b_{i}$ and $d_{i}$ are the birth and death times of a topological feature. To enrich this representation, we assign two additional geometric attributes for each point: (1) Persistence: $p_{i}=d_{i}-b_{i}$, measuring the lifespan of a feature; (2) Death-Birth Ratio: $\delta_{i}=log(\frac{d_{i}}{b_{i}})$, quantifying the relative scale of feature evolution. These are concatenated with the original birth and death values to form a 4-dimensional feature vector $v_{i}=(b_{i},d_{i},p_{i},\delta_{i})$ for each PD point, which is fed into the Persistence Diagram Encoder.
Point Cloud Data.
To provide geometric context, the model directly takes the raw 3D point cloud $C=\{c_{j}\}_{j=1}^{M}\subset\mathbb{R}^{3}$ as input (If the input is a 2D point cloud, we embed it into 3D space by setting the $z$ coordinate to zero). This allows the Point Cloud Encoder to learn geometric features directly from the source data from which the PD was computed.
Auxiliary Features.
To provide global context about the overall characteristics of both the topological and geometric data, we compute a 14-dimensional vector of auxiliary features, providing a comprehensive summary of the data’s global properties. This vector is fed into the Multi-Modal Fusion module and consists of statistical properties derived from both the persistence diagram and the point cloud:
  • PD Statistics (5 features): Number of points in the PD, mean persistence, standard deviation of persistence, maximum persistence, and mean birth time. These statistics capture the global topological distribution of the diagram. Since “significance” is relative rather than absolute, the maximum persistence acts as a normalization reference, guiding the distinction of true features from noise in diagrams with varying scales. Meanwhile, the mean and standard deviation quantify the background noise level, allowing the model to adaptively detect features based on the overall complexity and clutter of the topological summary.
  • Point Cloud Statistics (3 features): Number of points in the point cloud, the mean of the standard deviations across each coordinate axis (average spatial spread), and the mean Euclidean norm of the points (average distance from the origin). These metrics provide critical geometric scale and sampling context. Because persistence values are derived from geometric distances, they scale linearly with the object’s physical size. The average spatial spread and Euclidean norm enable the model to learn scale-invariance, preventing it from mistaking large-scale noise for signal in physically large objects.
  • Bounding Box Features (3 features): Axis-aligned extents along $x$, $y$, and $z$, computed as the per-axis maxmaxminmin differences of the point cloud coordinates. These features characterize the macro-geometry and volumetric density of the input. By combining spatial extents with the point count, the network can implicitly estimate the global sampling density, which governs the filtration speed and the expected birth times of loops. They also describe overall size and anisotropy, providing geometric context to interpret persistence magnitudes across shapes of different scales.
  • Noise Level and Uniformity Estimation (3 features): Given that the noise level and uniformity of point clouds significantly affects PDs, we estimate these properties of the point cloud by employing several statistical methods: (1) $k$-Nearest Neighbor Distance Standard Deviation: Measures variability in local densities by computing the standard deviation of per-point mean distances to $k=10$ nearest neighbors. A larger standard deviation indicates greater local density variation in the point cloud. (2) PCA-based Noise Estimate: Fits PCA on the entire point cloud and uses the ratio of the smallest to the largest eigenvalue of the explained variance as a noise indicator (set to 0 if the largest eigenvalue is non-positive). If the point cloud contains a low level of noise, the minimum eigenvalue should be much smaller than the maximum eigenvalue, and this ratio will be smaller. (3) Density Variation Coefficient: Uses the KNN-distance-based density proxy $1/\bar{d}$ and reports its coefficient of variation (standard deviation divided by mean) across points. A larger coefficient of variation indicates a more uneven global density distribution of the point cloud.
IV-B Model Architecture
Persistence Diagram Encoder.
We encode each persistence diagram as a set of enhanced per-point descriptors and aggregate them using attention [36], finally obtaining $g_{\text{pd}}$. Given raw pairs $(b_{i},d_{i})$, we construct a 4D point feature $v_{i}=[b_{i},d_{i},p_{i},r_{i}]$ with $p_{i}=d_{i}-b_{i}$ and $r_{i}=\log\tfrac{d_{i}}{b_{i}}$. Numerical stabilization is applied to avoid overflow and degeneracy. Each $v_{i}$ is processed by a point-wise MLP with standard nonlinearities and normalization, followed by a multi-head self-attention layer to model relationships among points. This yields contextualized per-point features $F_{i}$, from which a global diagram feature is obtained via pooling.
Point Cloud Encoder.
To incorporate geometric context, we use a point cloud encoder operating directly on raw point clouds, inspired by [32]. The point branch employs a stack of $1\times 1$ Conv1d layers, followed by normalization and ReLU, to generate features for each point. Subsequently, a compact global branch refines these representations, which are then aggregated into a global geometric descriptor $g_{\text{pc}}$ through a max-pooling operation.
Multi-Modal Fusion.
We project $g_{\text{pd}}$ and $g_{\text{pc}}$ into a common latent space and concatenate them with auxiliary features $a$. A small MLP with normalization produces a unified global context $f$.
Significance Classifier.
For each persistence point, we concatenate its contextualized feature $F_{i}$ with the fused global context $f$ and feed the result to a per-point classifier MLP with normalization to predict two-class logits. This design couples local topological salience with global geometric and topological context.
IV-C Loss Function
We adopt a weighted Focal Loss to address the severe class imbalance between rare significant points and abundant noise points. For each PD point $i$, let $\mathbf{s}_{i}\in\mathbb{R}^{2}$ be the two-class logits, $\mathbf{p}_{i}=\operatorname{softmax}(\mathbf{s}_{i})$ the class probabilities with components $\mathbf{p}_{i}[0],\mathbf{p}_{i}[1]\in[0,1]$, and $y_{i}\in\{0,1\}$ the ground-truth label (1: significant, 0: noise). Denote by $p_{t,i}=\mathbf{p}_{i}[y_{i}]$ the probability assigned to the true class. The per-point focal term is $\ell_{i}=w_{y_{i}}\,\alpha\,\big(1-p_{t,i}\big)^{\gamma}\,\big(-\log p_{t,i}\big)$, where $w_{y_{i}}\in\mathbb{R}_{+}$ is the class weight of label $y_{i}$ (typically $w_{1}>w_{0}$ to upweight the positive class), $\alpha>0$ is a global scaling coefficient, and $\gamma\geq 0$ is the focusing parameter that down-weights easy examples (large $p_{t,i}$) and emphasizes hard ones. Since PDs are padded to a fixed length, we compute the loss only over valid points via a mask $M_{i}\in\{0,1\}$ and normalize by the number of valid points:
$\mathcal{L}=\frac{1}{\sum_{i}M_{i}}\sum_{i}M_{i}\,\ell_{i}.$
In implementation, class weights enter through weighted cross-entropy, and the focal modulator multiplies this base term; masked positions are excluded via per-point sample weights consistent with the padding scheme.
TABLE I: Performance comparison of our model (TUN) against baselines (Mean $\pm$ Std for machine learning methods). More details of implementation and hyperparameters of comparison models are shown in Appendix -A.
Planar shapes (250 samples) CAD models (500 samples)
Model F1 Acc Pre Rec F1 Acc Pre Rec
2-means 0.6455 0.9483 0.5213 0.8474 0.4478 0.9292 0.3153 0.7725
CS(0.5) 0.2505 0.9521 0.9615 0.1440 0.1239 0.9652 0.9762 0.0662
CS(0.9) 0.1825 0.9498 0.9655 0.1008 0.1077 0.9649 0.9725 0.0570
SVM $0.3520_{0.0277}$ $0.9561_{0.0012}$ $0.9802_{0.0017}$ $0.2148_{0.0208}$ $0.1488_{0.0094}$ $0.9619_{0.0002}$ $4363_{0.0142} 0.0$$0897_{0.0063}$
MLP $0.7536_{0.0382}$ $0.9776_{0.0027}$ $0.9677_{0.0074}$ $0.6187_{0.0535}$ $0.5743_{0.1101}$ $0.9776_{0.0044}$ $0.9534_{0.0396}$ $0.4159_{0.1075}$
[…]
TUN 0.99670.0028 0.99960.0003 0.99770.0026 0.99570.0033 0.99780.0017 0.99980.0001 0.99620.0028 0.99920.0007
TPMS (500 samples) Zeolite (250 samples)
Model F1 Acc Pre Rec F1 Acc Pre Rec
2-means 0.4765 0.8231 0.3498 0.7474 0.5866 0.7501 0.4278 0.9327
CS(0.5) 0.0000 0.8923 0.0000 0.0000 0.0000 0.8099 0.0000 0.0000
CS(0.9) 0.0000 0.8923 0.0000 0.0000 0.0000 0.8099 0.0000 0.0000
SVM $0.1212_{0.0261}$ $0.8969_{0.0009}$ $7383_{0.0202} 0.0$$663_{0.0161} 0.0$$0007_{0.0015}$ $0.8099_{0.0001}$ $1455_{0.3253} 0.0$$0003_{0.0008}$
MLP $0.7262_{0.0674}$ $0.9527_{0.0091}$ $0.9511_{0.0115}$ $0.5909_{0.0858}$ $0.6456_{0.2854}$ $0.9070_{0.0566}$ $0.9259_{0.0579}$ $0.5596_{0.3283}$
LightGBM $0.9892_{0.0079}$ $0.9976_{0.0017}$ $0.9831_{0.0157}$ $0.9955_{0.0003}$ $0.9948_{0.0057}$ $0.9980_{0.0022}$ $0.9924_{0.0114}$ $0.9974_{0.0004}$
TUN 0.99870.0007 0.99970.0002 0.99850.0009 0.99880.0008 0.99960.0004 0.99980.0002 0.99990.0001 0.99920.0008
TABLE II: Detailed performance metrics (F1-score, Accuracy, Precision, Recall) for feature ablations (Mean $\pm$ Std).
Planar shapes CAD models
Model F1 Acc Pre Rec F1 Acc Pre Rec
Ablation 1 0.98160.0144 0.99800.0016 0.98540.0118 0.97830.0267 0.96410.0391 0.99740.0029 0.97220.0442 0.95800.0548
Ablation 2 0.98900.0133 0.99880.0014 0.98700.0250 0.99120.0062 0.99520.0034 0.99960.0003 0.99150.0078 0.99670.0029
TUN 0.99670.0028 0.99960.0003 0.99770.0026 0.99570.0033 0.99780.0017 0.99980.0001 0.99620.0028 0.99920.0007
TPMS Zeolite
Model F1 Acc Pre Rec F1 Acc Pre Rec
Ablation 1 0.99610.0021 0.99910.0005 0.99540.0043 0.99680.0046 0.99440.0026 0.99790.0010 0.99020.0070 0.99860.0030
Ablation 2 0.99790.0011 0.99960.0003 0.99880.0011 0.99710.0024 0.99830.0024 0.99940.0009 0.99990.0002 0.99680.0048
TUN 0.99870.0007 0.99970.0002 0.99850.0009 0.99880.0008 0.99960.0004 0.99980.0002 0.99990.0001 0.99920.0008
Figure 6: Illustrating a failure of TUN model. In this 1-PD of TPMS test sample, three points corresponding to small-scale loops in the original model have small persistence (marked in red in Ground Truth, while blue in Prediction), which makes them hard to distinguish from noise points.
V Results
In this section, we illustrate our experimental results. All experiments were conducted on a server equipped with one NVIDIA GeForce RTX 4090 GPU (24GB VRAM). The system was powered by a 16-core Intel Xeon Platinum 8352V CPU @ 2.10 GHz and utilized 120GB of system RAM.
V-A Datasets
We evaluate on four categories of point-cloud data: (i) planar shapes; (ii) CAD models drawn from the ABC dataset  [^1] https://deep-geometry.github.io/abc-dataset [24] and Thingi10k dataset  [^2] https://github.com/thingi10k/thingi10k [45]; (iii) triply periodic minimal surface (TPMS) structures  [^3] https://github.com/W-W-M/Morphology-Learning [37]; and (iv) zeolite crystal frameworks  [^4] http://america.iza-structure.org/IZA-SC/ftc_table.php [1]. Our dataset comprises a total of 5,000 samples, divided into training, validation, and test sets. The training and validation sets include 3,000 samples and 500 samples, respectively. They are allocated across all four classes. The test set contains 1,500 samples divided as follows: 250 planar shape samples, 500 CAD model samples, 500 TPMS samples, and 250 zeolite samples. Each point cloud is sampled from a shape whose first Betti number is fewer than 100, guaranteeing fewer than 100 significant points in the corresponding 1-PD. Point counts per point cloud range from hundreds to 100,000, ensuring diversity in density and noise level. Each PD sample is either truncated or padded to consist of 100 points, and each point cloud is subsampled or padded to contain 50,000 points.
For each instance, we construct a point cloud by sufficiently sampling the underlying geometric model and, when specified, perturbing points with additive noise at different levels. Some examples are illustrated in Figure 7 in Appendix -B. We then compute its 1D persistence diagram. Following the established identification rules for significant points, which align with the fundamental principles of topological data analysis, significance labels are assigned by category: for planar shapes, we annotate significant points via identifying and counting loops in the planar shapes; for 3D models whose sampled surfaces form 2D manifolds (closed surfaces), we compute the genus $g$, derive the 1D Betti number $\beta_{1}$ ($\beta_{1}=2g$ for closed orientable surfaces), and label the diagram’s significant birth-death pairs in accordance with $\beta_{1}$; for surfaces with boundary or multiple connected components, we compute Betti numbers by evaluating homology group ranks based on homology theory [16] and use these to guide the labeling; and if $\beta_{1}=0$, all points are labeled as noise. Owing to sufficient sampling, significant points typically emerge at early filtration times, exhibit large persistence, and their count matches the first Betti number $\beta_{1}$ of the underlying geometric model.
Crucially, our approach avoids circularity through a distinct information asymmetry between label generation and model inference. While ground-truth labels are derived using the exact $\beta_{1}$ from the underlying mesh, the TUN model never receives $\beta_{1}$ as input. Since $\beta_{1}$ varies dynamically across samples, the model cannot simply fit a fixed persistence threshold or rely on summary statistics (which means top-$k$ persistence selection for a fixed $k$ is unfeasible for handling all samples); instead, to correctly identify the specific set of significant points, TUN is forced to leverage the geometric features from the point cloud encoder to implicitly infer the topological complexity of the shape.
V-B Evaluation Metrics
We evaluate the performance of our model using standard classification metrics, including accuracy (Acc), precision (Pre), recall (Rec), and F1-score. The required TP, TN, FP, FN metrics for computing these metrics are given in Appendix -C. The best model is selected based on the lowest loss achieved on a held-out validation set, and the main evaluation metric for testing is F1-score. This choice is motivated by the severe class imbalance in our data: for each PD, we retain the top-100 points with the largest persistence, among which the number of positive labels (significant points) is typically much smaller than that of negative labels (noise points). F1-score, as the harmonic mean of precision and recall, provides a balanced measure that is especially sensitive to the model’s ability to correctly identify the minority positive class, making it the most informative metric for this imbalanced classification task.
V-C Results and Analysis
We present the quantitative results of our TUN in Table I. The evaluation is conducted on four distinct datasets. We first compare with two classic methods: the 2-means clustering method [17] and the Confidence Set (CS) method [14]. The 2-means clustering method partitions points in the 1-PD into two clusters—significant points and noise. The confidence-region method sets a persistence threshold and labels points above it as significant. In our experiments, we used the GUDHI toolbox  [^5] https://github.com/GUDHI/TDA-tutorial/blob/master/Tuto-GUDHI-ConfRegions-PersDiag-datapoints.ipynb  [27] to compute the confidence set for each 1-PD, testing confidence levels of 0.5 and 0.9 (denoted CS(0.5) and CS(0.9)) while keeping other parameters at their defaults. Next, we compare our TUN model with three machine-learning baselines: Support Vector Machine (SVM) [19], Multilayer Perceptron (MLP) [28], and LightGBM [23]. Every machine-learning method, including ours, is run 5 times; we report the mean and standard deviation of each evaluation metric. Additionally, we also compare our TUN full model against ablation variants that rely solely on persistence diagram features or exclude the auxiliary features.
Analysis and Comparison.
Our full model, TUN, demonstrates exceptional and state-of-the-art performance across all four datasets. As shown in Table I, it consistently achieves near-perfect F1-scores with low variance across all categories. Specifically, on the four datasets, TUN achieves F1-scores of $0.9967\pm 0.0028$, $0.9978\pm 0.0017$, $0.9987\pm 0.0007$, and $0.9996\pm 0.0004$ respectively. These results establish the superiority and robustness of our deep learning-based approach, confirming that TUN can consistently and accurately identify topological features in point clouds of varying complexity. Additionally, compared with traditional and standard supervised methods, the 2-means clustering yields poor F1-scores due to low precision, while the Confidence Set method is overly conservative, leading to near-zero recall. The SVM achieves low recall (0.00030.21480.0003-0.2148), failing to capture most significant features. The MLP achieves relatively high precision, but also suffers from moderate recall (0.410.620.41-0.62), failing to capture a significant portion of topological features. The LightGBM, a gradient boosting framework, achieves competitive F1-scores across the datasets. However, statistical aggregates struggle to capture intricate topological nuances in complex geometric structures. This limitation is evident on the CAD dataset, where LightGBM’s performance drops to an F1-score of $0.9688$ with a precision of $0.9521$. Such a precision gap implies a substantial number of false positive classifications, which can be unacceptable for high-stakes topological data analysis tasks. In contrast, TUN achieves superior classification results on the same dataset. This margin highlights the necessity of our multi-modal deep learning approach: while strong baselines like LightGBM reach a performance plateau, TUN successfully bridges the final gap to reliability by learning dependencies between geometry and topology that statistical summaries fail to represent.
Ablation Studies.
We conduct ablation studies to rigorously evaluate the contribution of each component in our multi-modal framework: (1) Ablation 1: Model only has the persistence diagram encoder, without the point cloud encoder and any auxiliary features. (2) Ablation 2: TUN model without all auxiliary features, only has persistence diagram encoder and point cloud encoder. Based on Table II, the results highlight the importance of multi-modal fusion. Ablation 1 yields respectable baselines but exhibits instability, particularly on the CAD dataset (0.9641) and Planar shapes (0.9816). This suggests that relying solely on persistence values is insufficient for robust classification, as it misses crucial spatial context. Ablation 2 incorporates the point cloud encoder, leading to substantial improvements. On CAD models, F1-score jumps to 0.9952, and on Zeolite, precision reaches near perfection (0.9999). This confirms that geometric features from the raw point cloud provide essential cues that help distinguish topological features from noise, especially in complex structures. Finally, the full TUN model, which further integrates auxiliary features, achieves the best overall performance. It attains the highest F1-scores across all datasets, demonstrating that both the point cloud encoder and the auxiliary features can enhance model stability and reliability.
V-D Limitations
In this study, we restrict the number of significant points in 1-PD to fewer than 100 and the number of point cloud points to fewer than 100,000. This is due to the consideration of computational efficiency based on the assumption that the significant points are no more than 100. If the number of points in the 1-PD and the point cloud is too large, the training process will be time-consuming and memory-intensive. But theoretically, our TUN model allows handling larger datasets with larger point clouds and PD with more points. Additionally, small-scale loops are difficult for TUN to detect reliably. For example, in Figure 6, the TUN model incorrectly predicts three points that resemble noise, which in fact correspond to three small-scale loops. One possible remedy is to adopt an extremely dense sampling strategy. However, ensuring the detection of such loops via dense sampling would incur substantial computational costs. Consequently, developing methods that can accurately and efficiently capture these features remains an open challenge.
VI Conclusion
We presented TUN, the first multi-modal neural network for significance detection in persistence diagrams that integrates topological, geometric, and auxiliary features. The method encodes diagram points with enhanced per-point descriptors, extracts global geometric context from raw point clouds, and fuses modalities through learned projections before per-point classification. This design bridges abstract topological features and their geometric origins, enabling accurate identification of truly significant features. Experimental results provide compelling evidence for the effectiveness of our proposed method, establishing a new state-of-the-art for topology understanding.
In future work, we will explore the topology understanding of higher-dimensional PDs, as well as explore alternative inputs instead of point clouds (like images, voxels) to broaden the applicability and improve the model’s interpretability.